OpenStudy (lchida25):

Calculate the bond enthalpy for the atomization of CCl4. I know how to calculate the actual bond enthalpy, but I'm not sure how to write the atomization of CCl4.

OpenStudy (jiteshmeghwal9):

In case of poly atomic molecules bond dissociation enthalpy is different for different bonds within the same molecule.

OpenStudy (jiteshmeghwal9):

\[CCl_4 \space \rightarrow \space C +4Cl \space \Delta _a H=?\]In such cases where all the bonds are identical in bond length & energy but the energy required to break the identical bond differs at each successive steps we use mean bond enthalpy. In \(CCl_4\), \[\Delta _{C-Cl}H =\frac{1}{4} \Delta _a H\]